OpenStudy (kpop4life123):

Jemma surveyed the students at her school to find out if they like the beach and/or the mountains for vacation. The table below shows the results of the survey: Like Beach Do Not Like Beach Total Like Mountains 36 28 64 Do Not Like Mountains 56 38 94 Total 92 66 158 If a student does not like the mountains, what is the probability that student also does not like the beach? (4 points) 39.1% 40.4% 59.5% 60.8%

OpenStudy (briannabug):

What do you think it is?

OpenStudy (kpop4life123):

b

OpenStudy (kpop4life123):

im not sure yet

OpenStudy (watermelon_talk):

lol

OpenStudy (kpop4life123):

omg lol

OpenStudy (briannabug):

Whats funny?

OpenStudy (watermelon_talk):

we both like kpop

OpenStudy (kpop4life123):

i need help guys :(

OpenStudy (briannabug):

To be honest I don't know how to explain it but if I were you I would most likely go with choice B. And what do you mean @Watermelon_Talk?

OpenStudy (watermelon_talk):

I think the answer is C

OpenStudy (briannabug):

It's either B or C

OpenStudy (kpop4life123):

no its b @briannabug got it right

OpenStudy (watermelon_talk):

total number of students who don like mountain+don like beach =56+38+94=188 total number of students=92+66+158=316 probability percentage =(total no. of favorable events/total number of events) x 100 =188/316 x 100 =59.49 ~ 59.5%

OpenStudy (watermelon_talk):

i got that from imqwerty

OpenStudy (kpop4life123):

but i got it right when i clicked b

OpenStudy (briannabug):

could someone help me?